# Arcade Processes for Informed Martingale Interpolation

### v2

@@ 2 Arcade processes
Suppose $K_{A}$ is of the form 2.18, we shall show that $(A_{t}^{(n)})$ is Markov. Let $k>1$ and $(s_{1},s_{2},\ldots,s_{k},t)\in[T_{0},T_{n}]^{k+1}$ such that $s_{1}<s_{2}<\ldots<s_{k}<t$. Then, $(A_{t}^{(n)})$ is Markov if and only if
$\mathbb{P}\left[A_{t}^{(n)}\in\cdot\mid A_{s_{1}}^{(n)},\ldots,A_{s_{k}}^{(n)}\right]=\mathbb{P}\left[A_{t}^{(n)}\in\cdot\mid A_{s_{k}}^{(n)}\right].$ (2.19)
Since $K_{A}(s,t)=0$ unless $s$ and $t$ are in the same sub-interval, we can assume without loss of generality that $(s_{1},s_{2},\ldots,s_{k},t)\in(T_{m},T_{m+1})^{k+1}$ for a fixed $m\in\{0,\ldots,n-1\}$. Define
@@
It remains to be shown that $A_{t}^{(n)}\perp\!\!\!\perp(\Delta_{1},\ldots,\Delta_{k-1})\text{ and }A_{s_{k}}^{(n)}\perp\!\!\!\perp(\Delta_{1},\ldots,\Delta_{k-1}).$ Equivalently, because we are treating the Gaussian case, we shall show that $\mathrm{Cov}[A_{t}^{(n)},\Delta_{q}]=0\mbox{ and }\mathrm{Cov}[A_{s_{k}}^{(n)},\Delta_{q}]=0\mbox{ for }q=1,\ldots,k-1.$ Expanding these covariances, we get
@@
The first equation implies the second because $K_{A}(s_{i},t)=0\implies K_{A}(s_{k},s_{i})=0$ by 2.18, and, when $K_{A}(s_{i},t)\neq 0$, we have $K_{A}(s_{i},t)=a(s_{k},t)K_{A}(s_{k},s_{i})$, so that
@@
For the converse, suppose without loss of generality that the driver has mean 00. We observe that $K_{A}(x,y)=0$ if $x\in[T_{i},T_{i+1}],y\in[T_{j},T_{j+1}],i\neq j$. Let $(r,s,t)\in(T_{i},T_{i+1})^{3}$ such that $r\leqslant s<t$. Since $(A_{t}^{(n)})$ is Gaussian, we have
$\mathbb{E}[A_{t}^{(n)}\,|\,A_{s}^{(n)}]=\frac{K_{A}(s,t)}{K_{A}(s,s)}A_{s}^{(n)}.$ (2.26)
@@
$K_{A}(r,t)=\mathbb{E}[A_{r}^{(n)}A_{t}^{(n)}]=\mathbb{E}[\mathbb{E}[A_{r}^{(n)}A_{t}^{(n)}\,|\,A_{s}^{(n)}]]=\mathbb{E}[\mathbb{E}[A_{r}^{(n)}\,|\,A_{s}^{(n)}]\mathbb{E}[A_{t}^{(n)}\,|\,A_{s}^{(n)}]],$ (2.27)
@@
A Gaussian AP $(A_{t}^{(n)})_{t\in[T_{0},T_{n}]}$ is Markov with respect to its own filtration if and only if there exist real functions $A_{1}:[T_{0},T_{n}]\rightarrow\mathbb{R}$ and $A_{2}:[T_{0},T_{n}]\rightarrow\mathbb{R}$ such that
@@
Conversly, suppose that $(A_{t}^{(n)})$ is Markov, and let $T_{m}\in\{T_{0},T_{n-1}\}_{*}$. Then,
@@
Let $T_{m}\in\{T_{0},T_{n-1}\}_{*}$ and $(s,t)\in(T_{m},T_{m+1})^{2}$ such that $s<t$. Choose the interpolating coefficients $f_{0},\ldots,f_{n}$ such that
@@
$K_{A}(x,y)=K_{D}(x,y)-\sum_{i=0}^{n}f_{i}(x)K_{D}(y,T_{i})=K_{D}(x,y)-\sum_{i=0}^{n}f_{i}(y)K_{D}(x,T_{i}),\quad\forall(x,y)\in[T_{0},T_{n}]^{2}.$ (2.38)
Let $T_{m}\in\{T_{0},T_{n-1}\}_{*}$ and $(s,t)\in(T_{m},T_{m+1})^{2}$ such that $s<t$. Recalling that $(D_{t})$ is Gauss-Markov, there exist two functions, $H_{1}:(T_{m},T_{m+1})\rightarrow\mathbb{R}$ and $H_{2}:(T_{m},T_{m+1})\rightarrow\mathbb{R}$, such that $K_{D}(s,t)=H_{1}(s)H_{2}(t)$. So, one can write
@@
Recall that we imposed on all drivers $(D_{t})$ the property that $\mathbb{P}[D_{t}=0]<1$ for $t\in(T_{0},T_{n})_{*}$. If we extend this property to $[T_{0},T_{n}]$ instead, then the above construction of a Markovian AP becomes explicit.
@@
If $(D_{t})$ is a Gauss-Markov process such that $\mathbb{P}[D_{t}=0]<1$ for $t\in[T_{0},T_{n}]$, where $K_{D}(s,t)=H_{1}(\min(s,t))H_{2}(\max(s,t))$ for all $(s,t)\in[T_{0},T_{n}]^{2}$ for some real functions $H_{1}$ and $H_{2}$, then the solution to Eq. (2.37) is given by
@@
Notice that if $(D_{t})$ does not satisfy $\mathbb{P}[D_{t}=0]<1$ for some $t\in\{T_{0},T_{n}\}_{*}$, it is still straightforward to construct the above AP by removing all the rows and columns of zeros in the matrix
@@
Since the driver already matches 00 at $t\in\{T_{0},T_{n}\}_{*}$, there is no need to find the corresponding interpolating coefficients because they do not appear in the AP expression. This is illustrated by taking the Brownian motion as a driver with $T_{0}=0$. Then $f_{0}$ does not matter since it is multiplied by 00 in the AP expression.
@@
This method to construct Markovian APs is not unique, but can be perceived as natural. The resulting APs are called standard.
@@
$A_{t}^{(n)}=\left\{\begin{array}[]{ll}D_{t}-\frac{H_{1}(T_{1})H_{2}(t)-H_{1}(t)H_{2}(T_{1})}{H_{1}(T_{1})H_{2}(T_{0})-H_{1}(T_{0})H_{2}(T_{1})}D_{T_{0}}-\frac{H_{1}(t)H_{2}(T_{0})-H_{1}(T_{0})H_{2}(t)}{H_{1}(T_{1})H_{2}(T_{0})-H_{1}(T_{0})H_{2}(T_{1})}D_{T_{1}}&\mbox{for }t\in[T_{0},T_{1}),\\ \\ D_{t}-\frac{H_{1}(T_{2})H_{2}(t)-H_{1}(t)H_{2}(T_{2})}{H_{1}(T_{2})H_{2}(T_{1})-H_{1}(T_{1})H_{2}(T_{2})}D_{T_{1}}-\frac{H_{1}(t)H_{2}(T_{1})-H_{1}(T_{1})H_{2}(t)}{H_{1}(T_{2})H_{2}(T_{1})-H_{1}(T_{1})H_{2}(T_{2})}D_{T_{2}}&\mbox{for }t\in[T_{1},T_{2}),\\ \,\vdots\\ D_{t}-\frac{H_{1}(T_{n})H_{2}(t)-H_{1}(t)H_{2}(T_{n})}{H_{1}(T_{n})H_{2}(T_{n-1})-H_{1}(T_{n-1})H_{2}(T_{n})}D_{T_{n-1}}&\\ \quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-\frac{H_{1}(t)H_{2}(T_{n-1})-H_{1}(T_{n-1})H_{2}(t)}{H_{1}(T_{n})H_{2}(T_{n-1})-H_{1}(T_{n-1})H_{2}(T_{n})}D_{T_{n}}&\mbox{for }t\in[T_{n-1},T_{n}],\end{array}\right.$ (2.50)
where $H_{1}$ and $H_{2}$ are the real functions appearing in the covariance function of the driver: $K_{D}(s,t)=H_{1}(\min(s,t))H_{2}(\max(s,t))$ for all $(s,t)\in[T_{0},T_{n}]^{2}$.
@@
$A_{t}^{(n)}=\left\{\begin{array}[]{l}D_{t}-\frac{\exp{[\theta(T_{1}-t)]}-\exp{[-\theta(T_{1}-t)]}}{\exp{[\theta(T_{1}-T_{0})]}-\exp{[-\theta(T_{1}-T_{0})]}}D_{T_{0}}-\frac{\exp{[\theta(t-T_{0})]}-\exp{[-\theta(t-T_{0})]}}{\exp{[\theta(T_{1}-T_{0})]}-\exp{[-\theta(T_{1}-T_{0})]}}D_{T_{1}}\quad\mbox{for }t\in[T_{0},T_{1}),\\ \\ D_{t}-\frac{\exp{[\theta(T_{2}-t)]}-\exp{[-\theta(T_{2}-t)]}}{\exp{[\theta(T_{2}-T_{1})]}-\exp{[-\theta(T_{2}-T_{1})]}}D_{T_{1}}-\frac{\exp{[\theta(t-T_{1})]}-\exp{[-\theta(t-T_{1})]}}{\exp{[\theta(T_{2}-T_{1})]}-\exp{[-\theta(T_{2}-T_{1})]}}D_{T_{2}}\quad\mbox{for }t\in[T_{1},T_{2}),\\ \,\vdots\\ D_{t}-\frac{\exp{[\theta(T_{n}-t)]}-\exp{[-\theta(T_{n}-t)]}}{\exp{[\theta(T_{n}-T_{n-1})]}-\exp{[-\theta(T_{n}-T_{n-1})]}}D_{T_{n-1}}-\frac{\exp{[\theta(t-T_{n-1})]}-\exp{[-\theta(t-T_{n-1})]}}{\exp{[\theta(T_{n}-T_{n-1})]}-\exp{[-\theta(T_{n}-T_{n-1})]}}D_{T_{n}}\\ \\ \hskip 324.36124pt\quad\mbox{for }t\in[T_{n-1},T_{n}].\end{array}\right.$ (2.52)
@@
If $(D_{t})$ is a standard Brownian motion, applying Eq. (2.37) to find appropriate interpolating coefficients yields the stitched Brownian AP, see Example 2.4. But there are other Markovian APs driven by standard Brownian motion. For instance, in the two-period case, we may choose
@@
$+(\min(s,t)-T_{1})\left(\frac{T_{0}(\max(s,t)+T_{0})-3T_{0}T_{1}+T_{1}^{2}}{(T_{1}-T_{0})^{2}}+\frac{T_{1}-\max(s,t)}{T_{2}-T_{1}}\right)\mathds{1}_{\left(T_{1},\frac{T_{1}+T_{2}}{2}\right]}(s,t)$
@@
×$\hskip 241.84842pt\times\mathds{1}_{\left(T_{1},\frac{T_{1}+T_{2}}{2}\right]}(\min(s,t))\mathds{1}_{\left(\frac{T_{1}+T_{2}}{2},T_{2}\right]}(\max(s,t))$
@@
×$\hskip 241.84842pt\times\mathds{1}_{\left(\frac{T_{1}+T_{2}}{2},T_{2}\right]}(s,t).$ (2.57)
@@
Recalling Theorem 3.8, this AP is Markovian. This process is a slight modification of the stitched Brownian AP, where $f_{0}$ is not 00 on $(T_{1},T_{2})$. Hence, $B_{T_{0}}$ still has a negative influence on the paths of the AP on $(T_{1},T_{2})$, since $f_{0}$ is negative on $(T_{1},T_{2})$. Similarly, we can modify the stitched Brownian arcade by setting $f_{2}\neq 0$ on $(T_{0},T_{1})$:
@@
$f_{1}(t)$ $=\frac{(t-T_{0})(T_{2}-T_{0})}{(T_{2}-T_{1})(T_{1}-T_{0})}\mathds{1}_{\left[T_{0},\frac{T_{0}+T_{1}}{2}\right]}(t)+\frac{T_{1}^{2}-T_{0}T_{2}+t(T_{0}-2T_{1}+T_{2})}{(T_{1}-T_{0})(T_{2}-T_{1})}\mathds{1}_{\left(\frac{T_{0}+T_{1}}{2},T_{1}\right]}(t)$
$+\frac{T_{1}(T_{0}(T_{1}-2T_{2})+T_{1}T_{2})+t(T_{0}T_{2}-T_{1}^{2})}{(T_{2}-T_{1})(T_{1}-T_{0})T_{1}}\mathds{1}_{\left(T_{1},\frac{T_{1}+T_{2}}{2}\right]}(t)$
$+\frac{(T_{2}-t)(T_{1}^{2}+T_{0}(T_{2}-2T_{1}))}{(T_{2}-T_{1})(T_{1}-T_{0})T_{1}}\mathds{1}_{\left(\frac{T_{1}+T_{2}}{2},T_{2}\right]}(t),$ (2.64)

### v3

@@ 2 Arcade processes
Suppose $K_{A}$ is of the form Eq. (2.12), we shall show that $(A_{t}^{(n)})$ is Markov. Let $k>1$ and $(s_{1},s_{2},\ldots,s_{k},t)\in[T_{0},T_{n}]^{k+1}$ such that $s_{1}<s_{2}<\ldots<s_{k}<t$. Then, $(A_{t}^{(n)})$ is Markov if and only if
$\mathbb{P}\left[A_{t}^{(n)}\in\cdot\mid A_{s_{1}}^{(n)},\ldots,A_{s_{k}}^{(n)}\right]=\mathbb{P}\left[A_{t}^{(n)}\in\cdot\mid A_{s_{k}}^{(n)}\right],$ (2.13)
see Theorem 1.12 in [20]. Since $K_{A}(s,t)=0$ unless $s$ and $t$ are in the same sub-interval, we can assume without loss of generality that $(s_{1},s_{2},\ldots,s_{k},t)\in(T_{m},T_{m+1})^{k+1}$ for a fixed $m\in\{0,\ldots,n-1\}$. Define
@@
It remains to be shown that $A_{t}^{(n)}\perp\!\!\!\perp(\Delta_{1},\ldots,\Delta_{k-1})\text{ and }A_{s_{k}}^{(n)}\perp\!\!\!\perp(\Delta_{1},\ldots,\Delta_{k-1}).$ Equivalently, because we are treating the Gaussian case, we shall show that $\mathrm{Cov}[A_{t}^{(n)},\\ \Delta_{q}]=0\mbox{ and }\mathrm{Cov}[A_{s_{k}}^{(n)},\Delta_{q}]=0\mbox{ for }q=1,\ldots,k-1.$ Expanding these covariances, we get
@@
The first equation implies the second because $K_{A}(s_{i},t)=0\implies K_{A}(s_{k},s_{i})=0$ by Eq. (2.12), and, when $K_{A}(s_{i},t)\neq 0$, we have $K_{A}(s_{i},t)=a(s_{k},t)K_{A}(s_{k},s_{i})$, so that
@@
For the converse, suppose without loss of generality that the driver has mean 0. We observe that $K_{A}(x,y)=0$ if $x\in[T_{i},T_{i+1}],y\in[T_{j},T_{j+1}],i\neq j$. Let $(r,s,t)\in(T_{i},T_{i+1})^{3}$ be such that $r\leqslant s<t$. Since $(A_{t}^{(n)})$ is Gaussian, we have
$\mathbb{E}[A_{t}^{(n)}\mid A_{s}^{(n)}]=\frac{K_{A}(s,t)}{K_{A}(s,s)}A_{s}^{(n)}.$ (2.20)
@@
$K_{A}(r,t)=\mathbb{E}[A_{r}^{(n)}A_{t}^{(n)}]=\mathbb{E}[\mathbb{E}[A_{r}^{(n)}A_{t}^{(n)}\mid A_{s}^{(n)}]]=\mathbb{E}[\mathbb{E}[A_{r}^{(n)}\mid A_{s}^{(n)}]\mathbb{E}[A_{t}^{(n)}\mid A_{s}^{(n)}]],$ (2.21)
@@
A Gaussian AP $(A_{t}^{(n)})_{t\in[T_{0},T_{n}]}$ is Markov with respect to its own filtration if and only if there exist real functions $A_{1}\mathrel{\mathop{\ordinarycolon}}[T_{0},T_{n}]\rightarrow\mathbb{R}$ and $A_{2}\mathrel{\mathop{\ordinarycolon}}[T_{0},T_{n}]\rightarrow\mathbb{R}$ such that
@@
Conversely, suppose that $(A_{t}^{(n)})$ is Markov, and let $T_{m}\in\{T_{0},T_{n-1}\}_{*}$. Then,
@@
Let $T_{m}\in\{T_{0},T_{n-1}\}_{*}$ and $(s,t)\in(T_{m},T_{m+1})^{2}$ be such that $s<t$. Recall Definition 2.1 and choose the interpolating coefficients $f_{0},\ldots,f_{n}$ so that they satisfy
@@
$K_{A}(x,y)=K_{D}(x,y)-\sum_{i=0}^{n}f_{i}(x)K_{D}(y,T_{i})=K_{D}(x,y)-\sum_{i=0}^{n}f_{i}(y)K_{D}(x,T_{i}),$ (2.32)
$\forall(x,y)\in[T_{0},T_{n}]^{2}$. Let $T_{m}\in\{T_{0},T_{n-1}\}_{*}$ and $(s,t)\in(T_{m},T_{m+1})^{2}$ such that $s<t$. Recalling that $(D_{t})$ is Gauss-Markov, there exist two functions, $H_{1}\mathrel{\mathop{\ordinarycolon}}[T_{m},T_{m+1}]\rightarrow\mathbb{R}$ and $H_{2}\mathrel{\mathop{\ordinarycolon}}[T_{m},T_{m+1}]\rightarrow\mathbb{R}$, such that $K_{D}(s,t)=H_{1}(s)H_{2}(t)$. So, one can write
@@
We recall that the property is imposed on all drivers $(D_{t})$ that for all $c\in\mathbb{R}$, $\mathbb{P}[D_{t}=c]<1$ for all $t\in(T_{0},T_{n})_{*}$. If we extend this property to $T_{0},T_{n}]$ instead, then the above construction of a Markovian AP becomes explicit.
@@
If $(D_{t})$ is a Gauss-Markov process such that for all $c\in\mathbb{R}$, $\mathbb{P}[D_{t}=c]<1$ for all for $t\in[T_{0},T_{n}]$, with $K_{D}(s,t)=H_{1}(\min(s,t))H_{2}(\max(s,t))$ for all $(s,t)\in[T_{0},T_{n}]^{2}$ for some real functions $H_{1}$ and $H_{2}$, then the solution to Eq. (2.31) is given by
@@
Notice that if there is a $c\in\mathbb{R}$ such that $\mathbb{P}[D_{t}=c]=1$ for some $t\in\{T_{0},T_{n}\}_{*}$, one can still use the construction of the above Markovian AP by removing all the rows and columns of zeros in the matrix
@@
This will determine all the interpolating coefficients but the ones that are associated with the times $T_{i}\in\{T_{0},T_{n}\}_{*}$ where $\mathbb{P}[D_{T_{i}}=c]=1$. Since the driver already matches $c$ at these times, their associated interpolating coefficients can be freely chosen without affecting the Markov property of the AP. This is illustrated by taking the Brownian motion with initial value $1$ as a driver with $T_{0}=0$. Then $f_{0}$ does not matter for the Markov property of the AP since it is multiplied by the deterministic function with constant value $1$ in the expression of the AP.
@@
This method to construct Markovian APs presented here is not unique, but is natural, for the covariance structure of the driver completely and uniquely determines the interpolating coefficients. The resulting APs are called standard. They are anticipative representations of processes whose law is determined by the conditioned driver, i.e., if $(A_{t}^{(n)})$ is standard, then $\mathbb{P}[A_{t}^{(n)}\in\cdot\,]=\mathbb{P}[D_{t}\in\cdot\mid D_{T_{0}}=0,D_{T_{1}}=0,\ldots,D_{T_{n}}=0]$.
@@
$A_{t}^{(n)}=\left\{\begin{array}[]{ll}D_{t}-\frac{H_{1}(T_{1})H_{2}(t)-H_{1}(t)H_{2}(T_{1})}{H_{1}(T_{1})H_{2}(T_{0})-H_{1}(T_{0})H_{2}(T_{1})}D_{T_{0}}\\ \\ \quad\quad\quad\quad\quad\quad\quad-\frac{H_{1}(t)H_{2}(T_{0})-H_{1}(T_{0})H_{2}(t)}{H_{1}(T_{1})H_{2}(T_{0})-H_{1}(T_{0})H_{2}(T_{1})}D_{T_{1}}&\mbox{for }t\in[T_{0},T_{1}),\\ \\ D_{t}-\frac{H_{1}(T_{2})H_{2}(t)-H_{1}(t)H_{2}(T_{2})}{H_{1}(T_{2})H_{2}(T_{1})-H_{1}(T_{1})H_{2}(T_{2})}D_{T_{1}}\\ \\ \quad\quad\quad\quad\quad\quad\quad-\frac{H_{1}(t)H_{2}(T_{1})-H_{1}(T_{1})H_{2}(t)}{H_{1}(T_{2})H_{2}(T_{1})-H_{1}(T_{1})H_{2}(T_{2})}D_{T_{2}}&\mbox{for }t\in[T_{1},T_{2}),\\ \\ \,\vdots\\ \\ D_{t}-\frac{H_{1}(T_{n})H_{2}(t)-H_{1}(t)H_{2}(T_{n})}{H_{1}(T_{n})H_{2}(T_{n-1})-H_{1}(T_{n-1})H_{2}(T_{n})}D_{T_{n-1}}&\\ \\ \quad\quad\quad\quad\quad\quad\quad-\frac{H_{1}(t)H_{2}(T_{n-1})-H_{1}(T_{n-1})H_{2}(t)}{H_{1}(T_{n})H_{2}(T_{n-1})-H_{1}(T_{n-1})H_{2}(T_{n})}D_{T_{n}}&\mbox{for }t\in[T_{n-1},T_{n}],\end{array}\right$. (2.44)
where $H_{1}$ and $H_{2}$ are the real functions that appear in the covariance function of the driver: $K_{D}(s,t)=H_{1}(\min(s,t))H_{2}(\max(s,t))$ for all $(s,t)\in[T_{0},T_{n}]^{2}$.
@@
$A_{t}^{(n)}=\left\{\begin{array}[]{ll}D_{t}-\frac{\exp{[\theta(T_{1}-t)]}-\exp{[-\theta(T_{1}-t)]}}{\exp{[\theta(T_{1}-T_{0})]}-\exp{[-\theta(T_{1}-T_{0})]}}D_{T_{0}}\\ \\ \quad\quad\quad\quad\quad\quad\quad-\frac{\exp{[\theta(t-T_{0})]}-\exp{[-\theta(t-T_{0})]}}{\exp{[\theta(T_{1}-T_{0})]}-\exp{[-\theta(T_{1}-T_{0})]}}D_{T_{1}}\quad&\mbox{for }t\in[T_{0},T_{1}),\\ \\ D_{t}-\frac{\exp{[\theta(T_{2}-t)]}-\exp{[-\theta(T_{2}-t)]}}{\exp{[\theta(T_{2}-T_{1})]}-\exp{[-\theta(T_{2}-T_{1})]}}D_{T_{1}}\\ \\ \quad\quad\quad\quad\quad\quad\quad-\frac{\exp{[\theta(t-T_{1})]}-\exp{[-\theta(t-T_{1})]}}{\exp{[\theta(T_{2}-T_{1})]}-\exp{[-\theta(T_{2}-T_{1})]}}D_{T_{2}}\quad&\mbox{for }t\in[T_{1},T_{2}),\\ \\ \,\vdots\\ \\ D_{t}-\frac{\exp{[\theta(T_{n}-t)]}-\exp{[-\theta(T_{n}-t)]}}{\exp{[\theta(T_{n}-T_{n-1})]}-\exp{[-\theta(T_{n}-T_{n-1})]}}D_{T_{n-1}}\\ \\ \quad\quad\quad\quad\quad\quad\quad-\frac{\exp{[\theta(t-T_{n-1})]}-\exp{[-\theta(t-T_{n-1})]}}{\exp{[\theta(T_{n}-T_{n-1})]}-\exp{[-\theta(T_{n}-T_{n-1})]}}D_{T_{n}}\quad&\mbox{for }t\in[T_{n-1},T_{n}].\end{array}\right$. (2.46)
@@
If $(D_{t})$ is a standard Brownian motion, applying Eq. (2.31) to find appropriate interpolating coefficients yields the stitched Brownian AP, see Example A.48. However, there are other Markovian APs driven by standard Brownian motion. For instance, in the two-period case, we may choose
@@
$+(\min(s,t)-T_{1})\left(\frac{T_{0}(\max(s,t)+T_{0})-3T_{0}T_{1}+T_{1}^{2}}{(T_{1}-T_{0})^{2}}+\frac{T_{1}-\max(s,t)}{T_{2}-T_{1}}\right)$
$\hskip 241.84842pt\times\mathds{1}_{\left(T_{1},\frac{T_{1}+T_{2}}{2}\right]}(s,t)$
@@
$\hskip 184.9429pt\times\mathds{1}_{\left(T_{1},\frac{T_{1}+T_{2}}{2}\right]}(\min(s,t))\mathds{1}_{\left(\frac{T_{1}+T_{2}}{2},T_{2}\right]}(\max(s,t))$
@@
$\hskip 241.84842pt\times\mathds{1}_{\left(\frac{T_{1}+T_{2}}{2},T_{2}\right]}(s,t).$ (2.51)
@@
Recalling Theorem 2.7, this AP is Markovian. This process is a slight modification of the stitched Brownian AP, where $f_{0}$ is not 0 on $(T_{1},T_{2})$. Hence, $B_{T_{0}}$ still has a negative influence on the paths of the AP on $(T_{1},T_{2})$, since $f_{0}$ is negative on $(T_{1},T_{2})$. Similarly, we can modify the stitched Brownian arcade by setting $f_{2}\neq 0$ on $(T_{0},T_{1})$:
@@
$f_{1}(t)$ $=\frac{(t-T_{0})(T_{2}-T_{0})}{(T_{2}-T_{1})(T_{1}-T_{0})}\mathds{1}_{\left[T_{0},\frac{T_{0}+T_{1}}{2}\right]}(t)$
$\quad+\frac{T_{1}^{2}-T_{0}T_{2}+t(T_{0}-2T_{1}+T_{2})}{(T_{1}-T_{0})(T_{2}-T_{1})}\mathds{1}_{\left(\frac{T_{0}+T_{1}}{2},T_{1}\right]}(t)$
$\quad+\frac{T_{1}(T_{0}(T_{1}-2T_{2})+T_{1}T_{2})+t(T_{0}T_{2}-T_{1}^{2})}{(T_{2}-T_{1})(T_{1}-T_{0})T_{1}}\mathds{1}_{\left(T_{1},\frac{T_{1}+T_{2}}{2}\right]}(t)$
$\quad+\frac{(T_{2}-t)(T_{1}^{2}+T_{0}(T_{2}-2T_{1}))}{(T_{2}-T_{1})(T_{1}-T_{0})T_{1}}\mathds{1}_{\left(\frac{T_{1}+T_{2}}{2},T_{2}\right]}(t),$ (2.58)